A spring has a natural length of 20cm. When a load of 50 N is applied, the spring becomes 22cm long in all. a) How great an extension does the 50 N force cause? b) If the spring obeys Hooke’s Law, what extension would you expect if a force of 150 N is applied to it? c) What will be the total length of the spring under that 150 N force?


1
Expert's answer
2019-09-23T09:19:07-0400

According to Hooke's law, "F = k \Delta x = k |l_1 - l_0|".

a) When the force "F_1 = 50 N" is applied, the length of the spring is "l_1 = 22 cm", so the extension is "\Delta_1 x= 22 cm - 20 cm = 2 cm".

b) From the first case, the spring constant according to Hooke's law is "k = \frac{F_1}{\Delta_1 x}" . The extension for the force "F_2 = 150 N" is "\Delta_2 x = \frac{F_2}{k} = \frac{F_1}{F_0} \Delta_1 x = 6 cm". The total length of the spring is "l_2 = l_0 + \Delta_2 x = 26 cm".